13. Factor the polynomial x^3-27. your answer can be written as (x-A)(x^2+bx+C)
7. If you rationalize the denominator of
1 / (11x sqrt(5) - 3y sqrt(3)
then you will get
A/B
what is A and B.
24.The difference of two positive numbers is 7 and the sum of their squares is 65. Find the numbers.

The bigger number is ,? and the smaller number is ?.
19.
Simplify the expression
[(4/x-1)-(1/x+1)] / [(x/x-1)+1/x+1)
and give your answer in the form of f(X)/ g(x)

Your answer for the function is f(x) :
Your answer for the function is g(x) :

21. The width of a rectangle is 7 less than twice its length. If the area of the rectangle is 69 cm, what is the length of the diagonal?

Note: Your answer must be a number. It may not contain any arithmetic operations.

The length of the diagonal is cm.


Answer by edjones(8007)

24)
x-y=7
x^2+y^2=65
.
x=y+7
(y+7)^2+y^2=65
y^2+14y+49+y^2=65
2y^2+14y-16=0
y^2+7y-8=0
(y+8)(y-1)=0
y=1, x=8
.
Ed